OpenStudy (anonymous):

For adults, the RDA of the amino acid lysine is 12 mg per kg of body weight. How many grams per day should a 79kg adult receive?

OpenStudy (anonymous):

12 * 79 = 948 mg 948/1000 = ?

OpenStudy (anonymous):

u mean this is the answer

OpenStudy (anonymous):

yes 1g = 1000mg

OpenStudy (anonymous):

ok what about this question <A typical multivitamin tablet can contain 2.0 mg of vitamin B2 (riboflavin), and the RDA is 0.0030 g/day. How many such tablets should a person take each day to get the proper amount of this vitamin, assuming that he gets none from any other sources?

OpenStudy (anonymous):

divide 0.0030 grams by 1000 to convert it in milligrams.

OpenStudy (anonymous):

sorry i meant multiply.

OpenStudy (anonymous):

then divide it by the amount of vitamin in a tablet.

OpenStudy (anonymous):

but he said in question Express your answer as an integer.

OpenStudy (anonymous):

You mean in decimal ?

OpenStudy (anonymous):

Like 2.3, 5.6, 1.2 etc.

OpenStudy (anonymous):

i dont know but he said after the question express your answer in integer

OpenStudy (anonymous):

what do you think about it

OpenStudy (anonymous):

ok that would be 3/2 = 1.5 ( one and a half tablet)

OpenStudy (anonymous):

0.0030 * 1000 = 3mg 3/2 = 1.5

OpenStudy (anonymous):

yes yes good result

OpenStudy (anonymous):

what about this question << RDA for the trace element selenium is 0.000070 g/day. Express this dose in mg/day.

OpenStudy (anonymous):

multiply it by 1000 and you have your answer.

OpenStudy (anonymous):

how can i do it in 2 significant figuers

OpenStudy (anonymous):

0.070 like this

OpenStudy (anonymous):

yes 0.07mg/day

OpenStudy (anonymous):

ok look about this question << While following a treasure map, you start at an old oak tree. You first walk 825 m directly south, then turn and walk 1.25 km at 30.0 ∘ west of north, and finally walk 1.00 km at 40.0 ∘ north of east, where you find the treasure: a biography of Isaac Newton!

OpenStudy (anonymous):

part1 <<< To return to the old oak tree, in what direction should you head ? Use components to solve this problem.

OpenStudy (anonymous):

What is your problem regarding this question ?

OpenStudy (anonymous):

While following a treasure map, you start at an old oak tree. You first walk 825 m directly south, then turn and walk 1.25 km at 30.0 west of north, and finally walk 1.00 km at 40.0 north of east, where you find the treasure: a biography of Isaac Newton!

OpenStudy (anonymous):

ok first convert all the identities in one mathematical system. For example convert all the kilometers in meters.

OpenStudy (anonymous):

What will you get ?

OpenStudy (anonymous):

Did you do it yet ?

OpenStudy (anonymous):

1 km = 1000 meters Multiply all the kilometers with 1000 to make them meters.